# Focused Angular $N$-Body Event Generator (FANG)

### v1

@@ Abstract
We introduce FANG (Focused Angular $n$-body event Generator), a new Monte Carlo tool for efficient event generation in restricted Lorentz-invariant phase space (LIPS). Unlike conventional approaches that uniformly sample the full $4\pi$ solid angle, FANG directly generates events in which selected final-state particles are constrained to fixed directions or finite angular regions in the laboratory frame. Because of the way the generator is constructed, angular constraints can be imposed directly in the laboratory frame while maintaining the correct LIPS structure, enabling differential and total cross sections or decay rates to be computed with high efficiency. The method is validated against analytic results and existing event generators, showing excellent agreement. By reducing computational cost by several orders of magnitude for angular observables, FANG provides a robust and versatile framework for applications in particle, nuclear, and detector physics.
I Introduction
Monte Carlo (MC) simulations are a standard tool in particle physics for generating events across the full LIPS [1, 2]. In these approaches, the squared matrix element is evaluated for each event, enabling the calculation of total cross sections. Differential cross sections, however, depend on the reference frame and must be evaluated in a specific frame. Of particular interest is the laboratory frame, where observables such as angular distributions are directly related to detector measurements. The methods developed here are general and can be extended to include both angular and energy differential cross sections in the laboratory frame.
When a detector is placed at a fixed angle, simulating events over the entire $4\pi$ solid angle is inefficient, since only a small portion of phase space contributes to the signal. Existing variance-reduction techniques [3, 4, 5, 6, 7, 8, 9, 10, 11] address this challenge in different contexts, but often require large event samples followed by cuts or interpolation to estimate detector yields. Because the probability of a particle arriving at exactly one angle is vanishingly small, interpolation or binning is necessary, which further increases computational cost.
In this work, we present FANG (Focused Angular $n$-body event Generator), a new MC algorithm that directly generates events with selected final-state particles constrained to specific directions or finite angular regions in the laboratory frame. The construction of the generator ensures exact kinematics and correct LIPS weighting while providing orders-of-magnitude improvements in efficiency for angular observables. This targeted approach eliminates the need for post-processing cuts and enables precise differential cross-section calculations in detector-relevant geometries.
II Theory
@@
where $P$ is the total four-momentum of the system, which is conserved during the interaction. The four-momentum of the $j$-th particle in the final state is denoted by $p_{j}$. The term $\delta(p_{j}^{2}-m_{j}^{2})$ enforces the on-shell condition for each particle, imposing $n$ constraints, where $m_{j}$ represents the mass of the $j$-th particle. $\delta^{4}\left(P-\sum_{j=1}^{n}p_{j}\right)$ enforces four-momentum conservation, providing 4 constraints. Together, these constraints define a (3n4)(3n-4)-dimensional manifold of the $4n$-dimensional space of all momentum components. Eq. 3 can be simplified to a sequential two-body decay problem, commonly referred to as the recurrence relation [1]. Introducing invariant masses $M_{i}$ and corresponding 4-momentum $q_{i}$ which satisfy $q_{i}^{2}=M_{i}^{2}$. We have:
@@
For the two-body phase space volume, one utilizes the property of Lorentz invariance and performs the calculation in the center-of-mass frame, resulting in:
@@
Here, the asterisk (*) denotes the two-body rest frame, where $\overset{\ast}{|\vec{p_{1}}|}={\sqrt{\overset{\ast}{E_{1}^{2}}-m_{1}^{2}}}$ and $\overset{\ast}{E_{1}}=\frac{M_{0}^{2}+m_{1}^{2}-m_{2}^{2}}{2M_{0}}$ are the momentum and energy of one of the particles in the two-body rest frame, and $d\Omega_{\overset{\ast}{p_{1}}}$ is the solid angle element of $\overset{\ast}{p_{1}}$ in the two-body rest frame. Note that Eq. 5 has 2 degrees of freedom, the azimuthal and polar coordinates, which aligns with the analysis made earlier, indicating that the system should have 3n43n-4 degrees of freedom. This is a special case for $n=2$. Alternatively, Eq. 5 may be written directly as a function of the particle masses [12]:
@@
II.1 Differential cross section in the laboratory frame
For the differential angular cross section of $n_{d}$ particles, with $n_{d}<n$, the particles are constrained to be emitted in fixed laboratory directions $(\hat{r}_{1},\hat{r}_{2},\dots,\hat{r}_{n_{d}})$. Referring to Eq. 8, in order to express all constrained solid angles in the laboratory frame, each factor $d\Omega_{\overset{\ast}{p_{k}}}$ must be replaced by
$J_{k}\,d\Omega_{p_{k}},\qquad k\in\{1,2,\dots,n_{d}\},$ (10)
where $J_{k}$ is the Jacobian relating the solid angle in the center-of-mass frame to that in the laboratory frame.
[…]
with $\theta_{k}$ and $\varphi_{k}$ denoting the polar and azimuthal angles of the momentum of particle $k$ in the laboratory frame, and $\overset{\ast}{\theta}_{k}$ and $\overset{\ast}{\varphi}_{k}$ the corresponding momentum angles in the center-of-mass frame, all angles defined relative to the two-body momentum direction. The Jacobian is evaluated at the angular coordinates $(\theta_{k},\varphi_{k})$ corresponding to the laboratory detector direction $\hat{r}_{k}$. For this pair of laboratory angles, the quantities $\overset{\ast}{\theta}_{k}$ and $\overset{\ast}{\varphi}_{k}$ are then determined according to the procedure described in Section II.2.
Computing this derivative yields [13]:
This expression for $J_{k}$ fails numerically at $\theta_{k}\to 0$, but the following limit is obtained [14]:
The $\gamma$ is the Lorentz factor of the two-body system in the lab frame, $\overset{*}{\beta}_{k}$ is the velocity of the $k^{th}$ constrained ejected particle in the two-body rest frame, and $\beta$ is the velocity of the two-body system in the lab frame.
Returning to Eq. 8, integrating this time only over solid angles of the unconstrained ejected particles and expressing the solid angle in the lab frame, the phase space element is expressed as
$\frac{d\sigma}{d\Omega_{1}\dots d\Omega_{k}}=$ $\frac{(2\pi)^{4-3n}}{4\sqrt{(P_{01}\cdot P_{02})^{2}-M^{2}_{01}M^{2}_{02}}}\frac{\pi^{n-1}}{2}\int_{M_{1}\dots M_{n-2}}|\mathcal{M}|^{2}F\left(\frac{m_{1}^{2}}{M_{0}^{2}},\frac{M_{1}^{2}}{M_{0}^{2}}\right)$ (16)
II.2 Calculating the four-momentum of the constrained particle
III Method
III.1 M-generation overview
FANG is based on the $M$-generation for n-body phase space generation, originally implemented in the code GENBOD [1]. The technique appears to have been developed independently by Kopylov [15] in the Soviet Union and by Raubold and Lynch at CERN. The $M$-generation algorithm treats the decay as a sequence of two-body decays (see Fig. 1), starting from the system mass $M_{0}$ and ending with the final two outgoing particles. For $n$ outgoing particles of masses $m_{i}$, there are n2n-2 intermediate virtual particles of masses $M_{i}$ that must be generated. The generation of the virtual masses $M_{i}$ proceeds according to the following prescription:
@@ III.1 M-generation overview
Starting from the $M_{0}$ rest frame, each two-body decay is treated by sampling $\cos\theta$ uniformly in the range [1,1][-1,1] and $\varphi$ uniformly in the range $0,2\pi]$ in the two-body center-of-mass frame, ensuring an isotropic phase space. The resulting four-momenta are then boosted back to the laboratory frame, after which the subsequent two-body decay can be evaluated. Analytically, the phase space for an $n$-body decay is obtained by integrating the differential volume element in Eq. 9 over the masses within the physical region. In the Monte Carlo approach, each event is assigned a weight equal to the integrand of Eq. 9. The phase space is then estimated by summing all weights corresponding to events in the physical region, multiplying by the volume spanned by the generated random numbers, and dividing by the total number of generated events. The range $\Delta M$ over which each virtual mass is generated is the same for all masses and is given by
@@
III.2 FANG
Building on on the $M$-generation described in Section (III.1), we extend the algorithm to allow, within an $n$-body phase space, laboratory-frame directional constraints on $n_{d}<n$ particles. This extension enables the generation of only those events consistent with the imposed constraints, while preserving exact kinematics and providing an accurate evaluation of the relative phase-space volume. For differential calculations, the constrained particle is generated in all events along the same laboratory-frame direction, specified by the vector $\hat{r}_{k}$. The kinematics of all other particles in the decay are, of course, influenced by this constraint. For partial phase-space event generation, where the constrained particle is restricted to a finite solid angle $\Delta\Omega_{k}$ in the laboratory frame, a new direction vector $\hat{r}_{k}$ is generated for each event, isotropically distributed within the limits of $\Delta\Omega_{k}$. This vector $\hat{r}_{k}$ is then used in the calculation of the four-momentum of the constrained particle. The inputs to the algorithm are the initial system four-momentum $P$, the outgoing particle masses $m_{i}$, and, for each constrained particle, the direction $\hat{r}_{k}$ together with the size and shape of the constraining solid angle $\Delta\Omega_{k}$. The algorithm then proceeds according to the following steps:
@@ III.2 FANG
III.2.1 Event weight calculation
@@
III.2.2 Differential calculation
The differential phase space in the specified laboratory-frame directions $\hat{r}_{k}$ is evaluated using point generation (possible solid-angle implementations include point, ring, strip, circular, and user-defined shapes; see Appendix A), where the constrained $k^{\text{th}}$ particle is directed along the same $\hat{r}_{k}$ for all events. The resulting differential phase space is given by
$\frac{dV_{n}(\hat{r}_{1},\dots,\hat{r}_{n_{d}})}{d\Omega_{1},\dots,d\Omega_{n_{d}}}=\frac{1}{N}\sum_{i=1}^{N}W_{i}$ (25)
@@ III.2.2 Differential calculation
$\frac{d\sigma(\hat{r}_{1},\dots,\hat{r}_{n_{d}})}{d\Omega_{1},\dots,d\Omega_{n_{d}}}=\frac{(2\pi)^{4-3n}}{4\sqrt{(P_{01}\cdot P_{02})^{2}-M^{2}_{01}M^{2}_{02}}}\frac{1}{N}\sum_{i=1}^{N}W_{i}|\mathcal{M}_{i}|^{2}$ (26)
@@
$\frac{dV_{n}(\theta_{1},\dots,\theta_{n_{d}})}{d\cos{\theta_{1}},\dots,d\cos{\theta_{n_{d}}}}=(2\pi)^{n_{d}}\frac{1}{N}\sum_{i=1}^{N}W_{i}$ (27)
@@
$\frac{d\sigma(\theta_{1},\dots,\theta_{n_{d}})}{d\cos{\theta_{1}},\dots,d\cos{\theta_{n_{d}}}}=\frac{(2\pi)^{4-3n}(2\pi)^{n_{d}}}{4\sqrt{(P_{01}\cdot P_{02})^{2}-M^{2}_{01}M^{2}_{02}}}\frac{1}{N}\sum_{i=1}^{N}W_{i}|\mathcal{M}_{i}|^{2}\,.$ (28)
@@
IV Results
IV.1 Full phase space calculation

### v2

@@ Abstract
We introduce FANG (Focused Angular $N$-body event Generator), a new Monte Carlo tool for efficient event generation in restricted Lorentz-Invariant Phase Space (LIPS). Unlike conventional approaches that sample the full $4\pi$ solid angle, FANG directly generates events in which selected final-state particles are constrained to fixed directions or finite angular regions in the laboratory frame. Because of the way the generator is constructed, angular constraints can be imposed directly in the laboratory frame while maintaining the correct LIPS structure, enabling differential and total cross sections or decay rates to be computed with high efficiency. The method is validated against analytic results and existing event generators, showing excellent agreement. By reducing the computational cost of full phase-space event generation by several orders of magnitude, FANG provides a robust and versatile framework applicable to particle, nuclear, and detector physics.
arxiv: 2509.11105
1 Introduction
The critical bridge between fundamental theory and experimental data is built upon the prediction of physical observables, such as cross sections or decay rates. These observables are computed by integrating the squared matrix element ($|\mathcal{M}|^{2}$) over the $n$-body Lorentz-Invariant Phase Space (LIPS). As the number of final-state particles $n$ increases, the dimensionality of this integration grows rapidly, rendering traditional numerical quadrature methods intractable. Consequently, Monte Carlo (MC) techniques are the only viable tools for performing this high-dimensional integration buckley_general-purpose_2011 . These MC techniques are implemented in software frameworks known as event generators, including Herwig bellm_herwig_2016 , Sherpa gleisberg_event_2009 , Pythia sjostrand_introduction_2015 , and MadGraph5_aMC@NLO alwall_automated_2014 , which are reviewed in buckley_monte_2019 ; valassi_challenges_2021 . In these generators one central task is the efficient sampling of the LIPS, as each sampled point represents a complete kinematic “event” of final-state particles. This kinematic sampling provides the kinematic configurations upon which the dynamics, governed by the integrand $|\mathcal{M}|^{2}$, are evaluated. While general-purpose algorithms capable of sampling the full LIPS exist such as RAMBO kleiss_new_1986 or GENBOD james_monte_1968 , their generality becomes a significant computational bottleneck when theoretical interest or experimental sensitivity is confined to a small sub-volume of the available phase space. This challenge arises, for example, from experimental fiducial cuts, which restrict measurements to a specific detector acceptance, as well as from dynamical effects, such as sharp resonances that concentrate the matrix element’s magnitude within very small regions of phase space bothmann_efficient_2023 ; kleiss_weight_1994 .
To overcome this inefficiency, standard MC techniques employ importance sampling strategies to concentrate event generation in regions where the integrand is largest. These methods, which include established adaptive multi-channel integration algorithms such as VEGAS lepage_adaptive_2021 and FOAM jadach_foam_2003 , as well as modern machine learning approaches such as generative neural networks (e.g., normalizing flows) butter_machine_2023 ; gao_event_2020 ; bothmann_exploring_2020 , adaptively build a sampling distribution that approximates the squared matrix element, $|\mathcal{M}|^{2}$. This allows for the preferential generation of events in dynamically favored regions. While highly effective, these strategies share a fundamental limitation when dealing with non-trivial kinematic constraints bothmann_exploring_2020 . While the physical integrand, $|\mathcal{M}|^{2}$, is itself Lorentz invariant, any sampling algorithm must parametrize the phase space using a specific set of coordinates to map random numbers to kinematic configurations. Event generation is almost always performed using a parametrization built from a chain of successive two-body decays (the so-called M-method), which is discussed in Section 2.1. The current samplers are thus meant to optimize the integration of the physics as a function of these two-body rest frame variables over their full and simple domains (e.g., $\cos\theta_{\mathrm{COM}}\in[-1,1]$). Experimental constraints, however, such as fiducial cuts or detector acceptance, are defined by simple geometric boundaries in the laboratory frame. Due to the non-trivial nature of the Lorentz boost, a region defined by simple, fixed-range constraints in the lab frame (e.g., $\theta_{\mathrm{lab}}\in[\theta_{1},\theta_{2}]$) maps to a complexly bounded, non-linear, and momentum-dependent sub-volume within the two-body rest frame parametrization. Standard generators, which efficiently sample the entire simple domain of the two-body rest frame coordinates, have no inherent knowledge of this disjoint target sub-volume. The only recourse is to apply these laboratory frame cuts a posteriori: one generates a large number of events across the full two-body systems domain, boosts them to the lab, and then discards the vast majority that fall outside the desired angular acceptance. This procedure is inefficient, with a computational cost that scales inversely with the acceptance volume, which can be prohibitively small.
In this paper, we present a novel approach that directly solves the problem of inefficient laboratory frame kinematic constraints. We introduce FANG (Focused Angular N-body event Generator), a new Monte Carlo framework designed specifically to generate $n$-body final states a priori within fixed, user-defined angular regions in the laboratory frame. The core of the FANG algorithm is a new phase space parametrization that incorporates these laboratory frame angular constraints directly into its construction. This allows the generator to sample only those kinematic configurations that satisfy the desired cuts, eliminating the need for wasteful a posteriori rejection. Critically, this is achieved while exactly preserving the Lorentz-invariant structure of the $n$-body phase space measure, ensuring that the generated event weights are correct. This technique enables the maximally efficient computation of differential and total cross sections or decay rates for observables defined in restricted angular acceptances, offering improvements reaching up to several orders of magnitude in computational efficiency over conventional methods.
This paper is organized as follows. Section 2 reviews the LIPS formalism and the recurrence relation underlying $n$-body kinematics, and summarizes the M-generation construction that we build upon. Section 3 introduces the theory and methodology of the FANG algorithm. Section 4 validates the method: (i) full–phase-space volumes against an algorithm implementing RAMBO; (ii) restricted generation against GENBOD with cuts; (iii) differential cross sections for $e\mu$ and $ep$ scattering benchmarked against analytic (Rosenbluth) results; and (iv) Demonstration of use with lepton–nucleus scattering models with comparison to GENIE andreopoulos_genie_2010 . Section 5 summarizes the results and outlines applications and future extensions.
2 Theoretical Background
@@
where $P$ is the total four-momentum of the system, which is conserved during the interaction. The four-momentum of the $j$-th particle in the final state is denoted by $p_{j}$. The term $\delta(p_{j}^{2}-m_{j}^{2})$ enforces the on-shell condition for each particle, imposing $n$ constraints, where $m_{j}$ represents the mass of the $j$-th particle. $\delta^{4}\left(P-\sum_{j=1}^{n}p_{j}\right)$ enforces four-momentum conservation, providing 4 constraints. Together, these constraints define a (3n4)(3n-4)-dimensional manifold of the $4n$-dimensional space of all momentum components. Eq. 3 can be simplified to a sequential two-body decay problem, commonly referred to as the recurrence relation james_monte_1968 . Introducing invariant masses $M_{i}$ and corresponding 4-momentum $q_{i}$ which satisfy $q_{i}^{2}=M_{i}^{2}$. We have:
@@
For the two-body phase space volume, one utilizes the property of Lorentz invariance and performs the calculation in the rest frame of the two bodies, resulting in:
@@
Here, the asterisk (*) denotes the two-body rest frame, where $\overset{\ast}{|\vec{p_{1}}|}={\sqrt{\overset{\ast}{E_{1}^{2}}-m_{1}^{2}}}$ and $\overset{\ast}{E_{1}}=\frac{M_{0}^{2}+m_{1}^{2}-m_{2}^{2}}{2M_{0}}$ are the momentum and energy of one of the particles in the two-body rest frame, and $d\Omega_{\overset{\ast}{p_{1}}}$ is the solid angle element of $\overset{\ast}{p_{1}}$ in the two-body rest frame. Note that Eq. 5 has 2 degrees of freedom, the azimuthal and polar coordinates, which aligns with the analysis made earlier, indicating that the system should have 3n43n-4 degrees of freedom. This is a special case for $n=2$. Alternatively, Eq. 5 may be written directly as a function of the particle masses kleiss_monte_2019 :
@@
2.1 M-generation overview
FANG is based on the $M$-generation for $n$-body phase space generation, originally implemented in the code GENBOD james_monte_1968 . The GENBOD technique appears to have been independently developed by Kopylov kopylov_notitle_1960 in the Soviet Union and by Raubold and Lynch at CERN. The $M$-generation algorithm treats the decay as a sequence of two-body decays (see Fig. 1), starting from the system mass $M_{0}$ and ending with the final two outgoing particles. For $n$ outgoing particles of masses $m_{i}$, there are n2n-2 intermediate virtual particles of masses $M_{i}$ that must be generated. The generation of the virtual masses $M_{i}$ proceeds according to the following prescription:
@@ 2.1 M-generation overview
Starting from the $M_{0}$ rest frame, each two-body decay is treated by sampling $\cos\theta$ uniformly in the range [1,1][-1,1] and $\varphi$ uniformly in the range $0,2\pi]$ in the two-body rest frame, ensuring an isotropic phase space. The resulting four-momenta are then boosted back to the laboratory frame, after which the subsequent two-body decay can be evaluated. Analytically, the phase space for an $n$-body decay is obtained by integrating the differential volume element in Eq. 9 over the masses within the physical region. In the Monte Carlo approach, each event is assigned a weight equal to the integrand of Eq. 9. The phase space is then estimated by summing all weights corresponding to events in the physical region, multiplying by the volume spanned by the generated random numbers, and dividing by the total number of generated events. The range $\Delta M$ over which each virtual mass is generated is the same for all masses and is given by
@@
3 FANG Approach
3.1 Differential cross section in the laboratory frame
To calculate the differential angular cross section of $n_{d}<n$ particles in fixed (constrained) $n_{d}$ directions in the laboratory frame, it is necessary to replace each factor of $d\Omega_{\overset{\ast}{p_{k}}}$ in Eq. 8 by
$d\Omega_{\overset{\ast}{p_{k}}}=J_{k}\,d\Omega_{p_{k}},\qquad k\in\{1,2,\dots,n_{d}\},$ (16)
where $J_{k}$ is the Jacobian relating the solid angle in the rest frame of the two bodies to that in the laboratory frame.
[…]
with $\theta_{k}$ and $\varphi_{k}$ denoting the polar and azimuthal angles of the momentum of particle $k$ in the laboratory frame, and $\overset{\ast}{\theta}_{k}$ and $\overset{\ast}{\varphi}_{k}$ the corresponding momentum angles in the two-body rest frame, all angles defined relative to this two-body system momentum direction. For each fixed (constrained) emitted laboratory directions $(\hat{r}_{1},\hat{r}_{2},\dots,\hat{r}_{n_{d}})$, the Jacobian is evaluated at the angular coordinates $(\theta_{k},\varphi_{k})$ corresponding to the laboratory direction $\hat{r}_{k}$. For this pair of laboratory angles, the quantities $\overset{\ast}{\theta}_{k}$ and $\overset{\ast}{\varphi}_{k}$ are then determined according to the procedure described in Section 3.2.
Computing this derivative yields horin_development_2025 :
This expression for $J_{k}$ fails numerically at $\theta_{k}\to 0$, but the following limit is obtained braithwaite_relativistic_1972 :
The $\gamma$ is the Lorentz factor of the two-body system in the laboratory frame, $\overset{*}{\beta}_{k}$ is the velocity of the $k^{th}$ constrained ejected particle in the two-body rest frame, and $\beta$ is the velocity of the two-body system in the laboratory frame.
Returning to Eq. 8, integrating this time only over solid angles of the unconstrained ejected particles and expressing the solid angle in the laboratory frame, the phase space element is expressed as
$\frac{d^{n_{d}}\sigma}{d\Omega_{1}\dots d\Omega_{n_{d}}}=$ $\frac{(2\pi)^{4-3n}}{4\sqrt{(P_{01}\cdot P_{02})^{2}-M^{2}_{01}M^{2}_{02}}}\frac{\pi^{n-1}}{2}\int_{M_{1}\dots M_{n-2}}|\mathcal{M}|^{2}F\left(\frac{m_{1}^{2}}{M_{0}^{2}},\frac{M_{1}^{2}}{M_{0}^{2}}\right)$ (22)
3.2 Calculating the four-momentum of the constrained particle
3.3 Methodology
Building on the $M$-generation described in Section 2.1, we extend the algorithm to allow, within an $n$-body phase space, laboratory-frame directional constraints on $n_{d}<n$ particles. This extension enables the generation of only those events consistent with the imposed constraints, while preserving exact kinematics and providing an accurate evaluation of the relative phase-space volume. For differential calculations, the constrained particle is generated in all events along the same laboratory-frame direction, specified by the vector $\hat{r}_{k}$. The kinematics of all other particles in the decay are, of course, influenced by this constraint. For partial phase-space event generation, where the constrained particle is restricted to a finite solid angle $\Delta\Omega_{k}$ in the laboratory frame, a new direction vector $\hat{r}_{k}$ is generated for each event, isotropically distributed within the limits of $\Delta\Omega_{k}$. This vector $\hat{r}_{k}$ is then used in the calculation of the four-momentum of the constrained particle. The inputs to the algorithm are the initial system four-momentum $P$, the outgoing particle masses $m_{i}$, and, for each constrained particle, the direction $\hat{r}_{k}$ together with the size and shape of the constraining solid angle $\Delta\Omega_{k}$. The algorithm then proceeds according to the following steps:
@@ 3.3 Methodology
3.3.1 Event weight calculation
@@
3.3.2 Integration over Detector geometry
For partial phase-space event generation, where a constrained particle is restricted to a finite solid angle $\Delta\Omega_{k}$ in the laboratory frame, a new constraining vector $\hat{r}_{k}$ is generated isotopically in the laboratory frame within $\Delta\Omega_{k}$ for each event. In these cases, the event phase space restricted to finite solid angles is given by the angle-integrated form of Eq. 13.
Where $W_{i}$ is the weight of each event given by Eq. 24.
For differential calculations, the algorithm supports the following routines:
3.3.3 Differential calculation
The differential phase space in the specified laboratory-frame directions $\hat{r}_{k}$ is evaluated using point generation, where the constrained $k^{\text{th}}$ particle is directed along the same $\hat{r}_{k}$ for all events. The resulting differential phase space is given by
$\frac{d^{n_{d}}V_{n}(\hat{r}_{1},\dots,\hat{r}_{n_{d}})}{d\Omega_{1},\dots,d\Omega_{n_{d}}}=\frac{1}{N}\sum_{i=1}^{N}W_{i}$ (26)
@@ 3.3.3 Differential calculation
$\frac{d^{n_{d}}\sigma(\hat{r}_{1},\dots,\hat{r}_{n_{d}})}{d\Omega_{1},\dots,d\Omega_{n_{d}}}=\frac{(2\pi)^{4-3n}}{4\sqrt{(P_{01}\cdot P_{02})^{2}-M^{2}_{01}M^{2}_{02}}}\frac{1}{N}\sum_{i=1}^{N}W_{i}|\mathcal{M}_{i}|^{2}$ (27)
@@
$\frac{d^{n_{d}}V_{n}(\theta_{1},\dots,\theta_{n_{d}})}{d\cos{\theta_{1}},\dots,d\cos{\theta_{n_{d}}}}=(2\pi)^{n_{d}}\frac{1}{N}\sum_{i=1}^{N}W_{i}$ (28)
@@
$\frac{d^{n_{d}}\sigma(\theta_{1},\dots,\theta_{n_{d}})}{d\cos{\theta_{1}},\dots,d\cos{\theta_{n_{d}}}}=\frac{(2\pi)^{4-3n}(2\pi)^{n_{d}}}{4\sqrt{(P_{01}\cdot P_{02})^{2}-M^{2}_{01}M^{2}_{02}}}\frac{1}{N}\sum_{i=1}^{N}W_{i}|\mathcal{M}_{i}|^{2}\,.$ (29)
@@
4 Results
4.1 Full phase space calculation